COMPARATIVE EXCITED STATE STUDIES OF SELECTED SPECIES USING QUANTUM MONTE CARLO

\underline{JOHN A. W. HARKLESS}, Physcial and Chemical Properties Division, National Institute of Standards and Technology, Gaithersburg, Maryland 20899-8380; RIAD MANAA, Lawrence Livermore National Laboratory, Livermore California 94550-9234; WILIAM A. LESTER, JR., Department of Chemistry, University of California, Berkeley, Berkeley, California 94720-0001 and Lawrence Berkeley National Laboratory, Berkeley, California 94720-0001.

Accurate estimates of excited state transition energies of selected molecular and atomic species are calculated using the quantum Monte Carlo method, other ab initio techniques, and density functional theory, where applicable. Some details of the different calculations are also discussed, in terms of not only computational expense, but also human expense.